# Automatic Forward Model Parameterization with Bayesian Inference of Conformational Populations

### v1

@@ II Theory
Second derivatives of the BICePs energy function and the BICePS score are useful for descent and uncertainty quantification using other forward models. We refrain from writing out the second derivative here, since the specific class of forward models we consider below all have second derivatives that go to zero. For more general cases, see Appendix A for more details. The energy of the Good-Bad likelihood model and its first and second derivatives are shown in Figure S1.
III Results/Discussion
Testing algorithm performance on a toy model
@@ III Results/Discussion
To investigate the efficacy of BICePs for this optimization problem, we introduce a simplified, yet comprehensive toy model. This model is designed to mimic the complexity of protein structure elements by generating $\phi$-angles from a multi-modal distribution, thereby emulating configurations characteristic of different secondary structure elements (Figure 1). This distribution encompasses three distinct modes, each characterized by a mean ($\mu$), standard deviation ($\sigma$), and weight ($w$): beta sheets ($\mu=-110^{\circ}$, $\sigma=20^{\circ}$, $w=0.35$), right-handed helices ($\mu=-60^{\circ}$, $\sigma=10^{\circ}$, $w=0.5$), and left-handed helices ($\mu=60^{\circ}$, $\sigma=5^{\circ}$, $w=0.15$). These parameters were chosen to accurately reflect the structural variability found in proteins. Angles $\phi_{i}$ were sampled from the multi-modal distribution,
@@
The sampled $\phi_{i}$ were then used to calculate experimental $J$-coupling constants $J(\phi)$ using the Karplus relation with the true Karplus coefficients ($A^{*}$, $B^{*}$, $C^{*}$).
${}^{3}\!J(\phi)=A\cos^{2}(\phi)+B\cos(\phi)+C$ (14)
Synthetic experimental $J$-coupling data is generated to represent a mixture of all conformational states, $d_{j}^{Exp}=\sum_{X}{{}^{3}\!J(\phi_{X,j})}\cdot p(X)$, with FM parameters $\theta=\{A,B,C\}$ set to their true values ($\{A=6.51,B=-1.76,C=1.6\}$). The initial forward model data is generated using reference Karplus parameters $(A_{0},B_{0},C_{0})$ and refined through the optimization process to showcase the algorithm’s adaptability and precision in parameter estimation.
Figure 1: A versatile toy model for measuring the performance of forward model optimization. The $\phi$-angles for each conformational state is pulled from a multi-modal distribution and corresponding energies. (a) This multi-modal distribution of $\phi$-angles was intended to represent configurations with different secondary structure elements having three distinct modes described by the mean ($\mu$), standard deviation ($\sigma$) and weight ($w$): beta sheets ($\mu=-110^{\circ}$, $\sigma=20^{\circ}$, $w=0.35$), right-handed helices ($\mu=-60^{\circ}$, $\sigma=10^{\circ}$, $w=0.5$), and left-handed helices ($\mu=60^{\circ}$, $\sigma=5^{\circ}$, $w=0.15$). (b) Cartoon representation of the backbone torsion angle, $\phi$.
BICePs robustly finds optimal Karplus parameters in the presence of experimental errors.
To evaluate the resilience of our algorithm against experimental inaccuracies, we introduced random and systematic errors of varying magnitudes ($\sigma_{\text{data}}$) into the synthetic experimental scalar couplings. The performance of our Good-Bad likelihood model, a Gaussian likelihood model, and singular value decomposition (SVD) was compared under these conditions.
@@
$\theta=V^{T}(\Sigma+\epsilon I)^{-1}U^{T}J^{\text{exp}},$ (16)
where $\epsilon=1e-6$ a small regularization term added to the diagonal of $\Sigma$ to ensure stability of the pseudo-inverse, and $J^{\text{exp}}$ represents the vector of experimental $J$-coupling values. This method ensures robust estimation of $\theta$ under ideal experimental conditions, given the true conformational populations. In practice, the true populations are not known a priori. The uncertainty in SVD coefficients was determined through 1k iterations of fitting, each omitting 10% of the data points chosen at random.
Typical uncertainties in NMR frequency measurements range from 0.1 to 1.0 Hz, primarily influenced by magnetic field strength, instrument quality, sample conditions, and the specifics of the pulse sequence used. In these experiments, 100 conformational states and 60 synthetic experimental scalar couplings were used. We introduced systematic error by shifting the experimental ${{}^{3}\!J}$ values by +2.0 Hz to +4.0 Hz for up to 20% of the data points. BICePs calculations were performed by averaging FM parameters over three chains of MCMC stating from different initial parameters ($\{A=9,B=-1,C=1\},\{A=4,B=0,C=3\},\{A=0,B=0,C=0\}$). Regardless of different starting parameters, posterior sampling universally converges to "true" optimal FM parameters. In these calculations, we used 32 BICePs replicas, and burned 10k steps followed by 50k steps of MCMC sampling.
@@
Our findings indicate that the Good-Bad likelihood model (red) exhibits superior resilience to experimental errors compared to a traditional Gaussian likelihood model (blue) and SVD (green) approaches (Figure 2). Predictions from SVD and the Gaussian likelihood model become notably less dependable when data incorporates errors, especially when $\sigma_{data}$ exceeds 0.5 Hz. On average, error in predictions (RMSE) from the Good-Bad model does not exceed 0.1 Hz over the full range of $\sigma_{data}$.
@@
Figure 2: Comparative analysis in performance of the Good-Bad likelihood model (red), a Gaussian likelihood model (blue), and singular value decomposition (SVD) using the "true" $\phi$ angles with synthetic experimental data. Here, we induced random and systematic error of varying magnitude ($\sigma_{data}$) to the experimental scalar couplings. Model performance was measured by computing RMSE (Hz) between the "true" scalar couplings and the couplings generated from the Karplus relations with predicted Karplus coefficients over 1,500 random perturbations to the experimental data, and represent the average of 100 BICePs calculations. Error bars represent the standard deviation. Predictions from SVD and the Gaussian likelihood model become notably less dependable when data incorporates errors, especially when $\sigma_{data}$ exceeds 0.5 Hz.
@@
An example of a single trial of forward model parameter refinement using the toy model is shown in Figure S2, where BICePs predicts Karplus coefficients by posterior sampling over FM parameters. Both BICePs and Singular Value Decomposition (SVD) methods successfully reproduce the "true" Karplus curve. However, BICePs excels by accurately identifying the error present in the data ($\sigma_{data}=0.471$), as indicated in the marginal posterior of uncertainty $p(\sigma_{J})$. The BICePs predicted maximum a posteriori uncertainty was found to be $\sigma_{J}=0.272$ with a variance scaling parameter of $\varphi_{J}=1.98$. The marginal posterior distributions of FM parameters for the Good-Bad model were $\{A=6.6\pm 0.04,B=-1.8\pm 0.02,C=1.5\pm 0.03\}$, and for SVD, $\{A=6.11\pm 0.06,B=-1.63\pm 0.04,C=1.80\pm 0.04\}$.
@@
In addition to the Good-bad model, we refined parameters using the Student’s model ($\{A=6.8\pm 0.03,B=-1.9\pm 0.03,C=1.4\pm 0.03\}$) to demonstrate that the Student’s model yields similar performance (Figure S3). The computed Gelman-Rubin ($\hat{R}$) statistic for these calculations was found to be $\hat{R}=1.01$ for each of the marginal posterior distributions of Karplus coefficients, which demonstrates that our chains converge to the same parameter location with similar variance.
@@
Furthermore, we assessed model performance across varying qualities of prior structural ensembles as illustrated in Figure S4. By introducing varying levels of prior error $\sigma_{prior}$ (measured in degrees) through perturbations to the "true" $\phi$ angles, even in the presence of random and systematic error, we observed strong correlation between the BICePs score and the quality of the structural ensemble, with a coefficient of determination $R^{2}$ of $0.99$. For these calculations, we employed the Good-Bad model, utilizing 32 replicas, and conducted 1,000 random perturbations to the $\phi$ angles with errors up to $\sigma_{prior}=4^{\circ}$, and perturbations to the experimental data $\sigma_{data}=0.68\pm 0.24$ Hz. Karplus’s warning about the perils of precise angle estimationKarplus (1963) underscores our approach’s necessity and performance in error aware modeling in structural biology.
@@
Variational minimization of the BICePs score to find optimal parameters.
Treating the forward model parameters as nuisance parameters, and sampling over them with the full posterior is an efficient strategy that grants the ability to include all sources of error while refining the structural ensemble with FM parameters. However, in the limit of large number of FM parameters, the dimensionality of the posterior may ultimately become unwieldy and present the curse of dimensionality. Here, we introduce an alternative strategy for refining FM parameters that has previously demonstrated to be a viable approach to automated force field optimizationRaddi and Voelz (2024).
In this approach, the FM parameters are no longer part of the joint posterior density. Instead, the posterior is conditioned on the set of FM parameters $\theta$, that is, equation 2 becomes
In this view, ensemble refinement is performed with a static set of FM parameters for each BICePs calculation.
$f(\theta)=-\ln\left(Z(\theta)\big{/}Z_{0}\right),$ (18)
where
[…]
is the evidence for FM parameters $\theta$, $Z_{0}$ is the evidence for a suitable reference state, and $u$ is the unchanged BICePs energy function (equation 9). To construct the reference state, we consider a series of likelihoods $p_{\xi}(D,\theta|\mathbf{X},\mathbf{\sigma})\sim[p(D|\mathbf{X},\mathbf{\sigma})]^{\xi}$ parameterized by $\xi\in[0,1]$, and set the reference state as the thermodynamic ensemble corresponding to $\xi=0$. The BICePs score is then calculated as the change in free energy of "turning on" experimental restraints ($\xi=0\rightarrow 1$).
It should be noted that in other applications of BICePs,Raddi et al. (2023); Raddi and Voelz (2024) the reference state for the BICePs score is defined using the $\lambda=0$ state for a series of a priors $p_{\lambda}(X)\sim[p(X)]^{\lambda}$, and the BICePs score is computed as the free energy of ($\lambda=0\rightarrow 1$) and ($\xi=0\rightarrow 1$) transformations. Here, since we are only interested in evaluating and/or parameterizing the likelihood functions, we set $p(X)$ to be uniform. Constructing $p(X)$ is thus very straightforward: it’s a collection of conformations all having equal statistical weight.
The derivative of the BICePs score with respect to the FM parameters $\theta$ reduces to the difference of Boltzmann averaged values of $\partial u/\partial\theta$ shown as
[…]
In this study, we demonstrate our methodology using first-order optimization methods, such as L-BFGS-B. For more complex forward models, the employment of second derivatives might become necessary. Interested readers are directed to the Supporting Information for second derivatives of the BICePs score with respect to FM parameters.
Calculation of the BICePs score (a free energy difference) and its derivatives (expectation values of energy derivative observables) is performed using the MBAR free energy estimator,Shirts and Chodera (2008) by sampling at several intermediates $\xi=0\rightarrow 1$, which enables accurate estimates of all quantities.
Optimizing $\xi$-values.
The accuracy of the BICePs scoredepends on converged sampling and sufficient thermodynamic overlap of intermediates ($\xi=0\rightarrow 1$) in the BICePs computation. To ensure strong overlap, we optimize the $\xi$-values by spacing ensembles equidistantly in thermodynamic length, employing a strategy akin to the "thermodynamic trailblazing" method proposed by Rizzi et al.Rizzi (2020) Our approach is facilitated by a custom optimization algorithm called pylambdaopt (Zhang et al., in preparation).
The optimization process is a two-step process: First, a preliminary BICePs calculation is performed using provisional $\xi$-values, yielding estimates of the thermodynamic length $|\ell(\xi_{n+1})-\ell(\xi_{n})|$ for each pair of intermediatesSivak and Crooks (2012); Shenfeld et al. (2009), derived from the variance in distributions $p(\Delta u_{n,n+1})$, where $\Delta u_{n,n+1}=u_{n+1}-u_{n}$ represents the change in the (reduced) BICePs energy incurred by bringing a sample from thermodynamic ensemble $n$ to thermodynamic ensemble $n+1$.
Second, cubic spline fitting is employed to derive a smooth and differentiable function $\ell(\xi)$ that accurately interpolates the computed $\ell(\xi_{i})$. Optimization through steepest-descent minimization is then applied to determine new $\xi_{i}^{*}$ values that minimize the loss function $\mathcal{L}=\sum_{n}|\ell(\xi_{n+1})-\ell(\xi_{n})|^{2}$. This results in $\xi_{i}^{*}$ values uniformly spaced in terms of thermodynamic length, thus maximizing the thermodynamic overlap between adjacent ensembles and enhancing the precision of free energy calculations. These optimized $\xi_{i}^{*}$ values are subsequently used in production runs. An illustration of the $\xi$-values pre- and post-optimization is depicted in Figure S5. Refer to figures S6&S7 for overlap matrices pre- and post- optimization.
@@
Comparison of variational minimization of the BICePs score vs. sampling the full joint posterior
@@
In the comparison of the two approaches for parameter estimation and optimization in our model, we utilized a toy model (Figure S2) to evaluate the efficacy of each method under the same data conditions. Prior to FM parameter refinement, 11 $\xi$-values were optimized from $\{1.0,0.9,0.8,...,0.0\}$ to $\{1.0,0.7,0.56,0.45,0.36,0.28,0.2,0.14,0.08,0.04,0.0\}$ (Figures S5-S7). Variational minimization using the Good-Bad model with 4 replicas (for reduced computational cost), where each evaluation of the objective function consisted of running 10k MCMC steps. Optimal parameters were determined to be $\{A=6.31\pm 0.02,B=-1.69\pm 0.03,C=1.69\pm 0.01\}$, averaged over 3 independent runs with very low variance between runs, shown in Figure S8. Regardless of different starting parameters ($\{A=9,B=-1,C=1\},\{A=4,B=0,C=3\},\{A=0,B=0,C=0\}$), variational minimization converges to “true” optimal FM parameters. This analysis demonstrated that both the joint posterior sampling approach and variational minimization yield near equivalent performance when applied to this model.
@@
The joint posterior sampling method, which involves sampling the joint posterior distribution of forward model (FM) parameters, has several advantages. One significant benefit is that the posterior distribution provides a direct estimate of the uncertainties in forward model parameters and their covariance. Compared to variational minimization, this method generally has a faster runtime and is particularly effective in handling non-convex landscapes, allowing for robust parameter estimation even in complex scenarios. However, it is not without drawbacks. As the number of FM parameters increases, the posterior sampling method may encounter the curse of dimensionality, which makes it computationally challenging to explore the parameter space efficiently.
@@
To evaluate the performance of our algorithm, we applied BICePs to human ubiquitin to predict Karplus coefficients for six sets of scalar coupling constants: ${{}^{3}\!J}_{H^{N}H^{\alpha}}$, ${{}^{3}\!J}_{H^{\alpha}C^{\prime}}$, ${{}^{3}\!J}_{H^{N}C^{\beta}}$, ${{}^{3}\!J}_{H^{N}C^{\prime}}$, ${{}^{3}\!J}_{C^{\prime}C^{\beta}}$, and ${{}^{3}\!J}_{C^{\prime}C^{\prime}}$. To test our algorithm’s robustness, we conducted a comprehensive evaluation for predicting optimal Karplus coefficients using three different structural ensembles as priors, each derived from distinct computational approaches: (1) 10 conformations from the NMR-refined structural ensemble, 1D3ZCornilescu et al. (1998), (2) 144 conformations from NMR-restrained simulations, 2NR2Richter et al. (2007), and (3) 25 conformations from the RosettaFold2 (RF2) algorithm.Baek et al. (2023)
We then validated the forward model parameters derived from each prior using the BICePs score, $R^{2}$ and mean absolute errors (MAE) for forward model predictions. As priors for these calculations, we used three independent structural ensembles: 1D3Z, 2NR2, and a 500-state conformational ensemble derived from a millisecond-long simulation of ubiquitin using CHARMM22*.Piana, Lindorff-Larsen, and Shaw (2013) For further details on these ensembles, refer to the SI methods section.
To refine the forward model (FM) parameters, we employed full joint posterior distribution sampling. This method was chosen to navigate the non-convex parameter space efficiently, given its relatively low dimensionality (18 FM parameters). BICePs calculations were executed by averaging the FM parameters over four Markov Chain Monte Carlo (MCMC) chains, each starting from distinct initial parameters: $\{A=9,B=-1,C=1\}$, $\{A=4,B=0,C=3\}$, $\{A=0,B=0,C=0\}$, and $\{A=6,B=-1,C=0\}$. Flexible residues were excluded from the calculations, consistent with previous studiesWang and Bax (1996); Habeck, Rieping, and Nilges (2005). As a result, a total of 346 $J$-couplings were used in these refinements. We used the Good-Bad model with 32 BICePs replicas, discarding the first 50k steps as burn-in, followed by 50k steps for MCMC sampling. Unlike the parameters derived from 1D3Z and RF2, the Karplus coefficients obtained by using the 2NR2 ensemble required a burn-in of 100k steps to appropriately converge due to a larger number of conformational states. The six sets of refined Karplus coefficients resulting from the 1D3Z, 2NR2 and RF2 ensembles are presented in Table 1.
@@
The predicted parameters, better represented by the marginal posterior distributions of the FM parameters, have large similarities across structural ensembles. BICePs-predicted coefficients using the 1D3Z ensemble (Figure S10) and predicted coefficients using the RF2 ensemble (Figure S11) are found to have very strong overlap. Furthermore, the traces of the FM parameters ober time (Figure S12) confirm convergence.
@@
One advantage of BICePs is that as FM parameters are being sampled, the posterior densities of FM uncertainties, $p(\sigma)$, are also revealed (Figure S13). For certain sets of $J$-coupling constants (e.g., ${{}^{3}\!J}_{H^{N}H^{\alpha}}$ and ${{}^{3}\!J}_{H^{N}C^{\prime}}$) the marginal posterior distribution of the variance scaling parameter $p(\varphi)$ has a sampled mean slightly larger than 1.0, indicating that the functional form of the likelihood opted for long tails to account for a few outlier data points deviating from the mean.

### v2

@@ II.1 Forward model optimization by posterior sampling of forward model parameters
Second derivatives of the BICePs energy function and the BICePs score are useful for descent and uncertainty quantification using other forward models. We refrain from writing out the second derivative here, but the derivation is straightforward. For the forward models we consider below (Karplus relations), all have second derivatives that go to zero. For more general cases, see Appendix Appendix A: Second partial derivatives of the BICePs score with respect to forward model parameters. for more details. The energy of the Good-Bad likelihood model and its first and second derivatives are shown in Figure S1.
II.2 Forward model optimization by variational minimization of the BICePs score
As we show in Results and Discussion, treating the forward model parameters as nuisance parameters and sampling them as part of the full posterior is an efficient strategy to refine FM parameters. However, in the limit of large numbers of FM parameters, the dimensionality of the posterior may become unwieldy and may suffer from the “curse of dimensionality”.
As an alternative strategy, we consider optimization of forward model parameters using variational minimization of the BICePs score, an approach that we have used to perform automated force field optimizationRaddi and Voelz (2024). In this approach, the FM parameters are no longer part of the joint posterior density. Instead, the posterior is conditioned on the set of FM parameters $\theta$, that is, equation 2 becomes
In this view, ensemble refinement is performed with a static set of FM parameters for each BICePs calculation. The BICePs score then enables the selection of the optimal model parameterized by $\theta$.
$f(\theta)=-\ln\left(Z(\theta)\big{/}Z_{0}\right),$ (15)
where
[…]
is the evidence for FM parameters $\theta$, $Z_{0}$ is the evidence for a suitable reference state, and $u$ is the unchanged BICePs energy function (equation 10). To construct the reference state, we consider a series of likelihoods $p_{\xi}(D,\theta|\mathbf{X},\mathbf{\sigma})\sim[p(D|\mathbf{X},\mathbf{\sigma})]^{\xi}$ parameterized by $\xi\in[0,1]$, and set the reference state as the thermodynamic ensemble corresponding to $\xi=0$. The BICePs score is then calculated as the free energy of "turning on" experimental restraints ($\xi=0\rightarrow 1$).
It should be noted that in other applications of BICePs,Raddi et al. (2025); Raddi and Voelz (2024) the reference state for the BICePs score is defined using the $\lambda=0$ state for a series of a priors $p_{\lambda}(X)\sim[p(X)]^{\lambda}$, and the BICePs score is computed as the free energy of ($\lambda=0\rightarrow 1$) and ($\xi=0\rightarrow 1$) transformations. Here, since we are only interested in evaluating and/or parameterizing the likelihood functions, we set $p(X)$ to be uniform.
To optimize forward model parameters, the BICePs score is used as an objective function that is estimated given a candidate value of $\theta$. Based on the estimate (and possibly other previous estimates), new candidates for $\theta$ are iteratively proposed and accepted or rejected, with the goal of finding the values $\theta^{*}$ that minimize the BICePs score.
For this purpose, gradient-based searches are very helpful. Conveniently, the derivative of the BICePs score $f(\theta)$ with respect to the FM parameters reduces to the Boltzmann-averaged values of the gradient of the $\partial u/\partial\theta$ :
[…]
In our results below, we use the first-order optimization method L-BFGS-B,Zhu et al. (1997) which only requires first derivatives. For more complex forward models, second-order methods may be needed. The Hessian matrix (of second partial derivatives of the BICePs score with respect to FM parameters) can be used for uncertainty estimation when $f(\theta)$ is convex. Assuming this is true, we estimate uncertainties in the FM model parameters using covariances obtained through inversion of the Hessian. Conceptually, this means that the estimated uncertainties in the best-fit values represent the widths of the basins in the BICePs score landscape. For non-convex problems, this can sometimes lead to an under-estimation of uncertainty depending upon the curvature of the basins of the local minima in the BICePs score landscape.
Calculation of the BICePs score (a free energy difference) and its derivatives (expectation values of energy derivatives) is performed using the MBAR free energy estimator,Shirts and Chodera (2008) by sampling at several intermediates $\xi=0\rightarrow 1$, which enables accurate estimates of all quantities. The accuracy of the BICePs score depends on converged sampling and sufficient thermodynamic overlap of intermediates ($\xi=0\rightarrow 1$) in the BICePs computation. To ensure strong overlap, we optimize the $\xi$-values by spacing ensembles equidistantly in thermodynamic length,Sivak and Crooks (2012); Shenfeld et al. (2009) employing a strategy akin to the "thermodynamic trailblazing" method proposed by Rizzi et al.Rizzi (2020) This optimization is performed using an algorithm we have developed called pylambdaopt.Novack et al. (2025)
III Results
A toy model to test the performance of forward model optimization
@@ III Results
To investigate the performance of BICePs in optimizing forward models, we introduce a simplified, yet physical, toy model. This model is designed to mimic the complexity of protein backbone structures by generating $\phi$-angles from a multi-modal distribution, thereby emulating configurations characteristic of different secondary structural elements (Figure 1). This distribution encompasses three distinct modes, each characterized by a mean ($\mu$), standard deviation ($\sigma$), and weight ($w$): beta sheets ($\mu=-110^{\circ}$, $\sigma=20^{\circ}$, $w=0.35$), right-handed helices ($\mu=-60^{\circ}$, $\sigma=10^{\circ}$, $w=0.5$), and left-handed helices ($\mu=60^{\circ}$, $\sigma=5^{\circ}$, $w=0.15$). These parameters were chosen to accurately reflect the structural variability found in proteins. Angles $\phi_{i}$ were sampled from the multi-modal distribution,
@@
The sampled $\phi_{i}$ were then used to calculate synthetic ensemble-averaged experimental $J$-coupling constants ${}^{3}J(\phi)$, using the Karplus relation
${}^{3}\!J(\phi)=A\cos^{2}(\phi)+B\cos(\phi)+C,$ (19)
with the coefficients set to their chosen true values: ($A^{*}$=6.51, $B^{*}$=-1.76, $C^{*}$=1.6). The synthetic $J$-coupling constant observables were calculated using the ensemble-average, as $\sum_{i}{{}^{3}J(\phi_{i}})p(\phi_{i})$.
Figure 1: A toy model for measuring the performance of forward model optimization. The $\phi$-angles for each conformational state is pulled from a multi-modal distribution and corresponding energies. (a) This multi-modal distribution of $\phi$-angles was intended to represent configurations with different secondary structure elements having three distinct modes described by the mean ($\mu$), standard deviation ($\sigma$) and weight ($w$): beta sheets ($\mu=-110^{\circ}$, $\sigma=20^{\circ}$, $w=0.35$), right-handed helices ($\mu=-60^{\circ}$, $\sigma=10^{\circ}$, $w=0.5$), and left-handed helices ($\mu=60^{\circ}$, $\sigma=5^{\circ}$, $w=0.15$). (b) Cartoon representation of the backbone torsion angle, $\phi$.
Including forward model parameters in the BICePs posterior is a robust method to optimize Karplus parameters in the presence of experimental errors.
Using the toy model described above, we next perform tests to examine whether including forward model parameters in the BICePs posterior can be used to find the optimal Karplus parameters starting from an initial reference forward model with parameters $(A_{0},B_{0},C_{0})$. Through these tests, we evaluate the precision of the results, and how robust the algorithm is in the presence of noise.
To evaluate the robustness of our algorithm in the presence experimental inaccuracies, we introduced random and systematic errors of varying magnitudes ($\sigma_{\text{data}}$) into the synthetic experimental scalar couplings. The performance of our Good-Bad likelihood model, a Gaussian likelihood model, and singular value decomposition (SVD) was compared under these conditions.
@@
$\theta=V(\Sigma+\epsilon I)^{-1}U^{T}J^{\text{exp}},$ (21)
where $\epsilon=1\times 10^{-6}$ is a small regularization term added to the diagonal of $\Sigma$ to ensure stability of the pseudo-inverse, and $J^{\text{exp}}$ represents the vector of experimental $J$-coupling values. This method ensures robust estimation of $\theta$ under ideal experimental conditions, given the true conformational populations. In practice, the true populations are not known a priori. The uncertainty in SVD coefficients was determined through 1000 iterations of fitting, each omitting 10% of the data points chosen at random.
Typical uncertainties in NMR frequency measurements range from 0.1 to 1.0 Hz, primarily influenced by magnetic field strength, instrument quality, sample conditions, and the specifics of the pulse sequence used. In these experiments, 100 conformational states and 60 synthetic experimental scalar couplings were used. We introduced systematic error by shifting the experimental ${{}^{3}\!J}$ values by +2.0 Hz to +4.0 Hz for up to 20% of the data points. BICePs calculations were performed by averaging FM parameters over three chains of MCMC starting from different initial parameters ($\{A=9,B=-1,C=1\},\{A=4,B=0,C=3\},\{A=0,B=0,C=0\}$). Regardless of different starting parameters, posterior sampling universally converges to "true" optimal FM parameters. In these calculations, we used 32 BICePs replicas, and burned 10k steps followed by 50k steps of MCMC sampling.
@@
Our findings indicate that the Good-Bad likelihood model (red) exhibits superior insensitivity to experimental errors compared to a traditional Gaussian likelihood model (blue) and SVD (green) approaches (Figure 2). Predictions from SVD and the Gaussian likelihood model become notably less dependable when the experimental data contains errors, especially when $\sigma_{\text{data}}$ exceeds 0.5 Hz. On average, error in predictions (RMSE) from the Good-Bad model does not exceed 0.1 Hz over the full range of $\sigma_{\text{data}}$.
@@
Figure 2: Comparative analysis in performance of the Good-Bad likelihood model (red), a Gaussian likelihood model (blue), and singular value decomposition (SVD) using the "true" $\phi$ angles with synthetic experimental data. Here, we induced random and systematic error of varying magnitude ($\sigma_{\text{data}}$) to the experimental scalar couplings. Model performance was measured by computing RMSE (Hz) between the "true" scalar couplings and the couplings generated from the Karplus relations with predicted Karplus coefficients over 1,500 random perturbations to the experimental data, and represent the average of 100 BICePs calculations. Error bars represent the standard deviation. Predictions from SVD and the Gaussian likelihood model become notably less dependable when data incorporates errors, especially when $\sigma_{\text{data}}$ exceeds 0.5 Hz.
@@
An example of a single trial of forward model parameter refinement using the toy model is shown in Figure S2, where BICePs predicts Karplus coefficients by posterior sampling over FM parameters. Both BICePs and SVD methods successfully reproduce the "true" Karplus curve. However, BICePs can accurately identify the error present in the data ($\sigma_{\text{data}}=0.471$ Hz), as indicated in the marginal posterior of uncertainty $p(\sigma_{J})$. The BICePs-predicted maximum a posteriori uncertainty was found to be $\sigma_{J}=0.272$ with a variance scaling parameter of $\varphi_{J}=1.98$. The marginal posterior distributions of FM parameters for the Good-Bad model were $\{A=6.6\pm 0.04,B=-1.8\pm 0.02,C=1.5\pm 0.03\}$, and for SVD, $\{A=6.11\pm 0.06,B=-1.63\pm 0.04,C=1.80\pm 0.04\}$.
@@
In addition to the Good-bad model, we refined parameters using the Student’s model ($\{A=6.8\pm 0.03,B=-1.9\pm 0.03,C=1.4\pm 0.03\}$) to demonstrate that the Student’s model yields similar performance (Figure S3). The computed Gelman-Rubin ($\hat{R}$) statisticGelman and Rubin (1992) for these calculations was found to be $\hat{R}=1.01$ for each of the marginal posterior distributions of Karplus coefficients, which demonstrates that our chains converge to the same parameter location with similar variance.
@@
Furthermore, we assessed model performance across varying qualities of prior structural ensembles as illustrated in Figure S4. By introducing varying levels of prior error $\sigma_{\text{prior}}$ (measured in degrees) through perturbations to the "true" $\phi$ angles, even in the presence of random and systematic error, we observed strong correlation between the BICePs score and the quality of the structural ensemble, with a coefficient of determination $R^{2}$ of $0.99$. For these calculations, we employed the Good-Bad model, utilizing 32 replicas, and conducted 1,000 random perturbations to the $\phi$ angles with errors up to $\sigma_{\text{prior}}=4^{\circ}$, and perturbations to the experimental data $\sigma_{\text{data}}=0.68\pm 0.24$ Hz.
@@
Variational minimization of the BICePs score to optimize FM parameters and comparison with posterior sampling
Next, we performed variational minimization of the BICePs score to determine optimal FM parameters for our toy model (Figure S2) , and compared the results to the posterior sampling approach under the same conditions.
@@
Variational minimization was performed using the Good-Bad model with 4 replicas (for reduced computational cost), where each evaluation of the objective function consisted of running 10,000 MCMC steps. Optimal parameters were determined to be $\{A=6.31\pm 0.02,B=-1.69\pm 0.03,C=1.69\pm 0.01\}$, averaged over 3 independent runs with very low variance between runs, shown in Figure S6. Regardless of different starting parameters ($\{A=9,B=-1,C=1\},\{A=4,B=0,C=3\},\{A=0,B=0,C=0\}$), variational minimization converges to “true” optimal FM parameters. This analysis demonstrated that both the joint posterior sampling approach and variational minimization show nearly equivalent performance when applied to this model.
@@
The posterior sampling method, which involves sampling the joint posterior distribution of forward model (FM) parameters, has several advantages. One significant benefit is that the posterior distribution provides a direct estimate of the uncertainties of the parameters and their covariance. Compared to variational minimization, this method generally has a faster runtime and is particularly effective in handling non-convex landscapes, allowing for robust parameter estimation even in complex scenarios. However, it is not without drawbacks. As the number of FM parameters increases, the posterior sampling method may encounter the curse of dimensionality, which makes it computationally challenging to explore the parameter space efficiently.
@@
As an application of our algorithm, we applied BICePs to human ubiquitin to predict Karplus coefficients for six sets of scalar coupling constants: ${{}^{3}\!J}_{H^{N}H^{\alpha}}$, ${{}^{3}\!J}_{H^{\alpha}C^{\prime}}$, ${{}^{3}\!J}_{H^{N}C^{\beta}}$, ${{}^{3}\!J}_{H^{N}C^{\prime}}$, ${{}^{3}\!J}_{C^{\prime}C^{\beta}}$, and ${{}^{3}\!J}_{C^{\prime}C^{\prime}}$. To test the robustness of our algorithm, we conducted a comprehensive evaluation using three different structural ensembles as priors, each derived from distinct computational approaches: (1) 10 conformations from the NMR-refined structural ensemble, 1D3ZCornilescu et al. (1998), (2) 144 conformations from NMR-restrained simulations, 2NR2Richter et al. (2007), and (3) 25 conformations from the RosettaFold2 (RF2) algorithm.Baek et al. (2023) Further details about these ubiquitin structural ensembles are given in Supporting Information.
We validated the forward model parameters derived from each prior using the BICePs score, $R^{2}$ and mean absolute errors (MAE) for forward model predictions. As priors for these validation calculations, we used three independent structural ensembles: 1D3Z and 2NR2 (described above), and a 500-state conformational ensemble derived from a millisecond-long simulation of ubiquitin using CHARMM22*.Piana, Lindorff-Larsen, and Shaw (2013) Further details about the CHARMM22* ensemble is given in Supporting Information.
To refine the forward model (FM) parameters, we employed full joint posterior distribution sampling. This method was chosen to navigate the non-convex parameter space efficiently, given its relatively low dimensionality (18 FM parameters). Estimates of FM parameters were made by averaging the BICePs results over four Markov Chain Monte Carlo (MCMC) sampling runs, each starting from distinct initial parameters: $\{A=9,B=-1,C=1\}$, $\{A=4,B=0,C=3\}$, $\{A=0,B=0,C=0\}$, and $\{A=6,B=-1,C=0\}$. Flexible residues were excluded from the calculations, consistent with previous studiesWang and Bax (1996); Habeck, Rieping, and Nilges (2005). As a result, a total of 346 $J$-couplings were used in these refinements. We used the Good-Bad model with 32 BICePs replicas, discarding the first 50k steps as burn-in, followed by 50k steps for MCMC sampling. Unlike the parameters derived from 1D3Z and RF2, the Karplus coefficients obtained by using the 2NR2 ensemble required a burn-in of 100k steps to appropriately converge due to a larger number of conformational states. The six sets of refined Karplus coefficients resulting from the 1D3Z, 2NR2 and RF2 ensembles are presented in Table 1.
@@
The predicted parameters have large similarities across structural ensembles. BICePs-predicted coefficients using the 1D3Z ensemble (Figure S8) and predicted coefficients using the RF2 ensemble (Figure S9) are found to strongly overlap. Furthermore, the traces of the FM parameters over time (Figure S10) confirm convergence.
@@
A key advantage of BICePs is its ability to sample the posterior densities of forward model uncertainties, $p(\sigma)$ (Figure S11). For certain sets of $J$-coupling constants (e.g., ${{}^{3}\!J}_{H^{N}H^{\alpha}}$ and ${{}^{3}\!J}_{H^{N}C^{\prime}}$) the marginal posterior distribution of the variance scaling parameter $p(\varphi)$ has a sampled mean slightly larger than 1.0, indicating that the functional form of the likelihood opted for long tails to account for a few outlier data points deviating from the mean.